there are 10 points in a plane of which 4 are collinear.
how many diff. straight lines can be drawn by joining these points..
(please explain the answer by using permutation combination)

Grade:10

7 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear ajinkya

straight line can be formed by joining 2 points

so number of ways in which we can select 2 points from 10 points is  = 10C2

 but it also include that in which 4 points are in straight line . from those 4 points only one line can be formed .

so we have to subtract 4C2 -1 from above result

so total lines are   = 10C2   - 4C2 + 1

 Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and

detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

Avijit Arya
14 years ago
Hi, For two points, one line can be drawn. For 3 points, 3 different lines can be drawn, on the condition that these points are not collinear, otherwise only single line will be drawn. Generalizing this, we can conclude that, for n number of points, we can draw a max. of nC2 lines. In this Q., there are 10 points, of which max. of 10C2, or 45 lines can be drawn. However, since 4 of these points are collinear, we have to subtract, the total no. of lines formed by these points, if they were not collinear, as that case is also included in the max. 45 lines formed. So subtracting 4C2 from 10C2 gives you =45-6=39 lines. Here we have subtracted all possible lines formed by those 4 points, however 1 line is formed by those 4 points, since they are collinear, so only 1 line is formed. So finally we have to add 1 line to the total we have got, that is 39. Hence the answer here is, 10C2-4C2+1=45-6+1= 40 lines, is the answer. Thanks
keerthi
9 years ago
from four points how only one straight can be formed
keerthi
9 years ago
from four points how only one straight can be formed
Hemanth
7 years ago
How many different straight lines can be formed by joining 12 different points on a plane of which four are collinear and the rest are non collinear?
Raghav bagla
6 years ago
I don`t have an answer I just wanted to ask that why can there be only 1 line formed with 4 points ( sorry for writing it here I didn`t know where to post my doubts )
ASh
6 years ago
25. There are 10 points in a plane, so straight lines can be formed using two at a time. \ Total straight lines = 10C2 = 10 9 2 1 × × = 45 1 But 4 points are collinear. \ Lines using these points = 4 C2 = 4 3 2 1 × × = 6 1 But these 4 points can make a single line. So, Total different lines = 45 – 6 + 1 = 40 1 Now, triangle can be formed using 3 points at a time. So, number of triangles = 10C3 = 10 9 8 321 × × × × = 120 1 Using 4 points taking 3 at a time. No. of triangles = 4 C3 = 4 1 But collinear points can’t make only tirangle. So, Total no. of triangles = 120 – 4 = 116